A vehicle is moving on circular track whose surface is inclined towards the horizon at an angle of 10degree the maximum velocity with which it can move safely is 36km/hr calculate the length of circular track

Grade:12

2 Answers

Vikas TU
6 years ago
Dears Student,
tan⁡∅=v2/rg
=>tan 10 = 10x10/rx10
=>r=56.71m
length= 2πr = 356.34m
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)
Shaikh Muzzakir
6 years ago
V= 36km/hr = 36×5÷18=10m/sAngle(@)=10°Formula: v=√rgtan@Squaring on both sides;we get v×v=rgtan@ 10×10=r×9.8×tan(10) 100=r×(9.8×0.1763) 100 =r(1.72774) r=100÷1.72774 r=57.8893~57.89m .....(1)We know that ;Length of the circular track(L) = 2πr L=2×3.142×57.89 ....from (1) (π=3.142) L=363.7808 m Muzzakir. (12th HSC Student Mumba)